# Small Bottle, Big Pipe: Quantifying and Addressing the Impact of Data Centers on Public Water Systems

### 03 Mar 2026 v1

@@ 1 Introduction
The high capital expenditure required to expand water capacity to meet the additional demand from data centers—as reflected by three recent water infrastructure upgrades for large technology companies costing a total of nearly \$1 billion  [16, 27, 18]—is frequently beyond the capability of many public water systems, particularly small and medium-sized systems which are often under-resourced and together constitute nearly 99% of community water infrastructure in the United States [28]. Like power grids, public water systems are designed to safely and reliably meet the maximum demand across all pressure zones, with margins to handle extreme conditions such as prolonged heatwaves and droughts. However, fulfilling this mission is becoming increasingly challenging due to longstanding underinvestment and limited technical resources. According to the EPA surveys on drinking water and wastewater systems [28, 29] and a recent third-party analysis [30], U.S. public water and wastewater systems now face a total infrastructure funding need of \$1.3–3.4 trillion over the next 20 years to improve service reliability, safety, and resilience. Moreover, as highlighted in the EPA’s report to Congress in 2024 [31], water affordability has become a significant challenge nationwide, with an estimate of 12.1 to 19.2 million U.S. households lacking affordable access to water services. These factors can further constrain the ability of public water systems to expand capacity for new industry-scale demands.
@@
To address the water capacity shortages without increasing burdens on local ratepayers, corporate-community partnerships are emerging to fund infrastructure upgrades and expansion. For example, a large technology company has recently committed up to \$400 million to invest in local public water infrastructure to support its new Louisiana data centers hosting AI and cloud services [16]. In another case, a water capacity request of 1.2 MGD from a data center developer (first reportedly for cooling [32] and later allegedly for fire suppression [33]) triggered major water infrastructure upgrades costing the developer over \$100 million [34]. Without these upgrades, the host town in Wisconsin, which had less than 2 MGD of remaining water capacity according to the water utility’s 2024 annual report [35], would likely see operational safety margin significantly reduced for its water system.
@@
Even when financial resources are available, however, public water infrastructure projects can take multiple years to complete and may face further delays due to additional constraints, including water rights allocation and multi-agency approvals [24, 36, 37]. In a recent instance, an Indiana water project, partly financed by the state, to add 25 MGD of new water supply capacity and expand wastewater treatment capacity by 15 MGD for an economic development district is expected to take approximately six years to complete and cost more than \$1 billion, including \$560 million for new water supply [38], \$225 million for upgrades to existing infrastructure, and \$350 million for wastewater treatment expansion [18]. As a result, a leading technology company’ data center located in that district to host AI and core products may not be able to operate at full capacity until its allocated water capacity of 8 MGD becomes available in 2031 [18, 19]. In a different and more complex instance, expanding the water reservoir capacity of an Oregon town located on federal land would require federal-level review and approval beyond the town’s jurisdiction [37].
@@
Compared with power grid expansion, public water system upgrades are often equally complex and time-consuming, and in some cases even more so, as they are heavily constrained by source availability, local hydrology, water rights, regulatory oversight, and/or environmental considerations [39, 40]. Moreover, many communities rely on reservoirs naturally replenished by rainfall or snowpack, which are highly sensitive to drought conditions and further constrain the development of additional reliable water supplies [24]. In addition, unlike cross-state electricity transmission, transporting water over long distances is considerably more challenging and generally not a viable option.
@@
Public water systems are designed to safely and reliably meet maximum demand at all times and across all pressure zones, with additional margins and dedicated fire flow capacity to address extreme conditions such as prolonged heatwaves and droughts. Accordingly, drinking water treatment plants, pump stations, storage tanks and reservoirs, transmission mains, distribution networks, and wastewater treatment facilities must all be appropriately sized and carefully planned to accommodate peak scenarios. As a result, capacity expansion and upgrades to water infrastructure often require multiple years to complete, particularly for complex projects such as treatment plant expansions. These projects typically involve environmental review, water rights permitting and allocation, multi-agency approvals, funding authorization, detailed engineering design, and construction, all of which can substantially extend implementation timelines [24, 36].
@@ 2.3 Water Capacity Bottleneck for Data Centers
Unfortunately, many U.S. public water systems lack sufficient available capacity to accommodate new, large-scale industrial peak demands [24, 28]. Concretely, depending on local climate conditions and cooling system design, supporting a 100 MW IT load using evaporative cooling, which is modest relative to emerging gigawatt-scale AI facilities, can require a peak water capacity allocation of approximately 0.5–2.5 MGD (Appendix C). This estimate excludes deployments in hot climates, where elevated ambient temperatures can substantially increase water capacity need and push the required peak withdrawal capacity beyond 5 MGD. By comparison, a town in Wisconsin with roughly 12,000 residents has less than 2 MGD of remaining physical capacity, according to its 2024 annual report [35]. Thus, after accounting for operational safety margins and reliability headroom, the practically allocable surplus capacity is likely even smaller, which may be difficult to support the need of a evaporatively cooled 100 MW IT load, let alone data centers at larger scales. In fact, many data center projects have required substantial upgrades to local water infrastructure, even when their peak water demand is as low as approximately 0.1 MGD [75, 76, 77, 21, 18, 16, 34].
@@
To balance grid stress and water use, some data centers have increasingly adopted hybrid cooling strategies, such as applying evaporative cooling to a portion of IT loads and/or supplementing dry coolers with direct evaporative assistance only during summer periods. Nonetheless, even if only 10% of IT loads are cooled with direct evaporative assistance during the summer in a cold climate, the growing scale of AI data centers can still lead to substantial peak water demand. For example, when fully built, an AI-dedicated data center is planned for approximately 0.7 MGD of peak water demand, which can be substantial for the available surplus capacity of a small or medium public water system in the host community in Wisconsin [78]. In another case of a host town with less than 2 MGD available remaining physical capacity (according to the water utility’s 2024 annual report) [35], a peak water capacity request of 1.2 MGD reportedly triggers a major water infrastructure upgrade exceeding \$100 million [34]. More recently, a leading technology company’s state-of-the-art data center under construction employs a “closed-loop, liquid-cooled system that recirculates the same water” [19] and relies on water evaporation for facility-level cooling only during the hottest days of the year. However, the data center may not be able to operate at full capacity until its allocated water capacity of 8 MGD becomes available, as anticipated in 2031 [18, 19].
@@ 3.1.1 High consumptive ratio
Table 1: 2024 U.S. data center water and energy use of selected companies. An asterisk “” denotes values estimated or assigned based on third-party sources. A dagger “” denotes ratios derived from reported water consumption values when withdrawal-based ratios are not explicitly disclosed. A dash “–” indicates that the corresponding value is not reported and cannot be plausibly estimated. The details of data sources are available in Appendix A.4.
@@ 3.1.2 High peaking factor
Figure 2c presents the monthly water withdrawal of a hyperscale data center in West Des Moines, Iowa, indicating a ratio of peak monthly withdrawal to average monthly withdrawal of 4.30. It is important to note that the disclosed withdrawal volume includes a small portion of domestic water use, which typically exhibits a peaking factor in the range of 1.5 to 2.5. Consequently, the monthly peak-to-average water withdrawal ratio attributable solely to the data center cooling load is likely higher than 4.30. Applying a minimum adjustment factor of 1.5 to translate average demand to maximum day demand [83] yields an estimated daily peaking factor of at least 6.45.
We further present in Figure 2b the percentage contribution of the data center’s monthly water use relative to the total water use of the top 20 largest users in West Des Moines. This analysis further highlights the pronounced seasonal variability of data centers that rely on evaporative assistance during the summer.
@@
Figure 2: Water use data collected from the monthly financial reports of West Des Monies Water Works available at [48]. (a) Monthly water use of a hyperscale data center in West Des Moines, Iowa; (b) The ratio of the data center’s water use to the total of 20 largest water users served by West Des Moines Water Works; (c) Monthly peaking factors (i.e., ratio of the maximum monthly water use to the average use between 2022 and 2025) of all the large water users counted by West Des Moines Water Works.
Although the data center’s water use is relatively low during cooler months, its elevated summer water demand is sufficiently large that it becomes the largest annual water user in both 2024 and 2025 served by West Des Moines Water Works [48].
By comparison, we also present in Figure 2c the monthly peaking factors for other large water users reported in the financial disclosures of West Des Moines Water Works from 2022 to 2025 [48]. Some water users appear in the reports for fewer than four years; for these users, we compute the ratio of their maximum monthly water use to their average monthly use over the period in which they are reported. We can see that all other large water users exhibit substantially lower peaking factors than the data center. Note that, water user T, a large corporate campus, exhibits an abnormally high water use of approximately 15 million gallons in August 2023, exceeding more than ten times its average monthly use between 2022 and 2025. This spike is likely associated with campus expansion and infrastructure upgrades during that period, as indicated by the City of West Des Moines, Iowa building permit reports (May to July 2023) [86], resulting in a one time increase in water use. It does not exhibit the recurring annual pattern observed for the data center and other users. We therefore exclude this anomalous month when calculating the monthly peaking factor for user T.
@@
The water use of data centers exhibits pronounced seasonal and even daily spikes, resulting in peaking factors estimated to range from 3 to 10, depending on ambient temperature and cooling system configuration. Cooling systems that rely on evaporative assistance during the summer generally exhibit substantially higher peaking factors than conventional evaporative cooling towers. Additionally, in cooler climates and/or when server-level cooling operates at higher temperature setpoints that only need evaporative assistance for a limited number of days, the peaking factor can be even higher. For planning purposes, similar to power capacity redundancy [47, 88], data centers often request more water than their actual demand to hedge against extreme heat events and/or rapid future load increases. Finally, we note that although water storage tanks can buffer short-term fluctuations in hourly demand, data centers rarely rely on storage to fully offset multi-day peak water demand because of the high costs associated with utility-scale tanks, sanitation, maintenance, and other operational requirements.
@@ 4.1.2 Scenarios
While individual companies may achieve greater WUE reductions, the industry-wide adjustments applied in our study are consistent with previous observations and align with common reduction targets reported by data center operators [106]. Importantly, under the Moderate and Optimistic scenarios, the projected U.S.-wide average WUE in 2030 (Table 7) is approximately 0.43 L/kWh and 0.31 L/kWh, respectively. These values are lower than the modeled WUE projections reported in [2] and are consistent with prevailing industry trends, thereby representing plausible forward-looking estimates. In addition to the Baseline, Moderate, and Optimistic scenarios, we include two references—Reference (LBNL) and Reference (NS)—which apply WUE values based on the modeled results in [2] and [5], respectively. [^4] For our analysis, we re-calculate the total water consumption for Reference (LBNL) rather than directly using the 2024–2028 values reported in [2]. Specifically, we follow the standard annual on-site WUE definition, resulting in lower water consumption estimates than those originally reported. See Table 8 in the appendix for details.

### 18 Mar 2026 v2

@@ 1 Introduction
The high capital expenditure required to expand water capacity to meet the additional demand from data centers—as reflected by three recent water infrastructure upgrades for large technology companies costing a total of nearly \$1 billion  [16, 27, 18]—is frequently beyond the capability of many public water systems (Table 1), particularly small and medium-sized systems which are often under-resourced and together constitute nearly 99% of community water infrastructure in the United States [28]. Like power grids, public water systems are designed to safely and reliably meet the maximum demand across all pressure zones, with margins to handle extreme conditions such as prolonged heatwaves and droughts. However, fulfilling this mission is becoming increasingly challenging due to longstanding underinvestment and limited technical resources. According to the EPA surveys on drinking water and wastewater systems [28, 29] and a recent third-party analysis [30], U.S. public water and wastewater systems now face a total infrastructure funding need of \$1.3–3.4 trillion over the next 20 years to improve service reliability, safety, and resilience. Moreover, as highlighted in the EPA’s report to Congress in 2024 [31], water affordability has become a significant challenge nationwide, with an estimate of 12.1 to 19.2 million U.S. households lacking affordable access to water services. These factors can further constrain the ability of public water systems to expand capacity for new industry-scale demands.
@@
To address water capacity shortages without increasing the burden on local ratepayers, state laws often require large water users to make “pro rata” fair contributions to fund infrastructure upgrades and system expansion [32, 33], sometimes in combination with corporate-community partnerships. For example, a large technology company has recently committed up to \$400 million to invest in local public water infrastructure to support its new Louisiana data centers hosting AI and cloud services [16]. In another case, a water capacity request of 1.2 MGD from a data center developer (first reportedly for cooling [34] and later allegedly for fire suppression [35]) triggered major water infrastructure upgrades costing the developer over \$100 million [36]. Without these upgrades, the host town in Wisconsin, which had less than 2 MGD of remaining water capacity according to the water utility’s 2024 annual report [37], would likely see operational safety margin significantly reduced for its water system.
@@
Even when financial resources are available, however, public water infrastructure projects can take multiple years to complete and may face further delays due to additional constraints, including engineering water rights allocation and multi-agency approvals [24, 38, 39]. In a recent instance, an Indiana water project, partly financed by the state, to add 25 MGD of new water supply capacity and expand wastewater treatment capacity by 15 MGD for an economic development district is expected to take approximately six years to complete and cost more than \$1 billion, including \$560 million for new water supply [40], \$225 million for upgrades to existing infrastructure, and \$350 million for wastewater treatment expansion [18]. As a result, a leading technology company’ data center located in that district to host AI and core products may not be able to operate at full capacity until its allocated water capacity of 8 MGD becomes available in 2031 [18, 19].
In some cases, the availability of local and regional water sources can become a key and complex bottleneck. For example, expanding the capacity of a reservoir located on federal land for an Oregon town would require federal-level review and approval beyond the town’s jurisdiction [39], which can significantly constrain the water supply available to new large users or expansions of existing demand.
In another instance involving a data center project in Virginia, the county record [33] indicates that the future expansion of water capacity from 2 MGD to 8 MGD to serve the data center “will limit the water supply available [to the water authority’s service area] and will impair economic and other development opportunities in those localities that are dependent on water supply.” The county further recognizes that “new water supply sources must be identified, accessed, and developed” for the water and wastewater system [33]. This requirement effectively accelerates the water authority’s long-range water source planning by decades [41]. As a result, future expansion of the data center project may face substantial delays while new water supply sources are developed, a process which itself is a “lengthy and costly project” [33].
@@
Compared with power grid expansion, public water system upgrades are often equally complex and time-consuming, and in some cases even more so, as they are heavily constrained by source availability, local hydrology, water rights, regulatory oversight, and/or environmental considerations [42, 43]. For example, it can take roughly 20 years or more to search for a new water source and have it in service [32]. Moreover, many communities rely on reservoirs naturally replenished by rainfall or snowpack, which are highly sensitive to drought conditions and further constrain the development of additional reliable water supplies [24]. In addition, unlike cross-state electricity transmission, transporting water over long distances is considerably more challenging and generally not a viable option.
@@
Public water systems are designed to safely and reliably meet maximum demand at all times and across all pressure zones, with additional margins and dedicated fire flow capacity to address extreme conditions such as prolonged heatwaves and droughts. Accordingly, drinking water treatment plants, pump stations, storage tanks and reservoirs, transmission mains, distribution networks, and wastewater treatment facilities must all be appropriately sized and carefully planned to accommodate peak scenarios. As a result, capacity expansion and upgrades to water infrastructure often require multiple years to complete, particularly for complex projects such as treatment plant expansions. These projects typically involve environmental review, water rights permitting and allocation, multi-agency approvals, funding authorization, detailed engineering design, and construction, all of which can substantially extend implementation timelines [24, 38]. For example, the process of identifying, developing, and bringing a new water source into service can take roughly 20 years or longer [32].
@@ 2.3 Water Capacity Bottleneck for Data Centers
Table 1: U.S. water treatment surplus capacity (in MGD) by system service population category for all water sources and systems based on the EPA’s most recent nationwide community water system survey (2006) [78]. The average surplus is defined as the design capacity minus the peak daily flow (Table 19 in [79]). *As the median surplus is not reported, we use the median design capacity minus the median peak daily treatment production (Table 9 in [80], which does not further disaggregate systems serving “over 10,000” people by size).
Category
100 or
Less
101 -
500
501 -
3,300
3,301 -
10,000
10,001 -
50,000
50,001 -
100,000
100,001-
500,000
Over
500,000
All Sizes
Average Surplus 0.08 0.14 0.34 0.68 0.84 2.48 3.86 9.41 0.72
Median Surplus* 0.03 0.10 0.23 0.89 3.00 0.20
Unfortunately, many U.S. public water systems lack sufficient available capacity to accommodate new, large-scale industrial peak demands [24, 28]. Table 1 summarizes the surplus water treatment capacity of U.S. community water systems based on the EPA’s most recent nationwide survey conducted in 2006 [78]. The results indicate that the available surplus capacity is limited across all system sizes when compared with the water capacity demand of large data centers. Systems relying on surface water generally have larger total and surplus capacities than those relying on groundwater sources [79]. Although this survey has not been updated, the current situation is likely similar or potentially more constrained, as reflected by the rapidly growing water infrastructure funding needs (after inflation adjustment) shown in the EPA’s 2023 Report to Congress [28]. Meanwhile, water capacity expansions are typically implemented incrementally in line with population growth, making it unlikely that the current nationwide statistics of water treatment surplus capacity differs substantially from the values reported in Table 1. For example, under California’s Drinking Water State Revolving Fund (DWSRF) policies [81], funding for capacity expansions is generally limited to 10% above the existing maximum daily demand. This rule is intended to ensure engineering reliability, maintain service to existing customers, and reduce financial and operational risk.
Moreover, according to California’s 2025 report on water resources [82], more than 50% of urban water suppliers in California report zero annual surplus capacity, while an additional 4.5% report annual shortages that require supplier response actions. Among suppliers reporting monthly data, 20% project shortages during certain months of the coming year that require shortage response measures. Although the data on daily supply and demand is not available, the fraction of suppliers experiencing shortages of daily capacity for some days of the year is likely substantially higher than 20%. This suggests that accommodating new large water users, such as data centers that rely on evaporative cooling during the summer, can be challenging for many urban water suppliers under existing capacity constraints.
Table 1 reports only the treatment plant capacity, while other critical components such as distribution pipes, wastewater processing, storage, and water source availability can also be bottlenecks. In addition, after accounting for operational safety margins, reliability headroom, fire suppression requirements, usage categories, and distribution losses, the practically allocable surplus capacity becomes even smaller. The allocable capacity is the most relevant factor for connecting large water users, even when the system’s total capacity is large. For example, in Norwalk, Iowa, a city with approximately 15,000 residents, only 0.65 MGD of water capacity is designated as industrial reserve for new large water users as of March 2026, even though other existing water uses may be more substantial [83].
Consequently, depending on local climate conditions and cooling system design, supporting a 100 MW IT load using evaporative cooling—which is modest compared with emerging gigawatt-scale AI facilities and generally requires a peak water capacity allocation of approximately 0.5–2.5 MGD (Appendix C)—can frequently exceed the available surplus capacity of public water systems. This estimate does not include deployments in hot climates, where elevated ambient temperatures can substantially increase water demand, potentially pushing the required peak withdrawal capacity beyond 5 MGD. In fact, many data center projects have already required substantial upgrades to local water infrastructure, even when their water capacity demand is as low as approximately 0.1 MGD [84, 85, 86, 21, 18, 16, 36].
@@
To balance grid stress and water use, some data centers have increasingly adopted hybrid cooling strategies, such as applying evaporative cooling to a portion of IT loads and/or supplementing dry coolers with direct evaporative assistance only during summer periods. Nonetheless, even if only 10% of IT loads are cooled with direct evaporative assistance during the summer in a cold climate, the growing scale of AI data centers can still lead to substantial peak water demand. For example, when fully built, an AI-dedicated data center is planned for approximately 0.7 MGD of peak water demand, which can be substantial for the available surplus capacity of a small or medium public water system in the host community in Wisconsin [87]. In another case of a host town with less than 2 MGD available remaining physical capacity (according to the water utility’s 2024 annual report) [37], a peak water capacity request of 1.2 MGD reportedly triggers a major water infrastructure upgrade exceeding \$100 million [36]. More recently, a leading technology company’s state-of-the-art data center under construction employs a “closed-loop, liquid-cooled system that recirculates the same water” [19] and relies on water evaporation for facility-level cooling only during the hottest days of the year. However, the data center may not be able to operate at full capacity until its allocated water capacity of 8 MGD becomes available, as anticipated in 2031 [18, 19]. In another instance involving a new data center project in Virginia, the county record [33] indicates that the future expansion of water capacity from 2 MGD to 8 MGD to serve the data center “will limit the water supply available [to the water authority’s service area] and will impair economic and other development opportunities in those localities that are dependent on water supply.” As a result, future expansion of the data center project may face substantial delays while new water supply sources are developed, a process which itself is a “lengthy and costly project” [33].
@@ 3.1.1 High consumptive ratio
Table 2: 2024 U.S. data center water and energy use of selected companies. An asterisk “” denotes values estimated or assigned based on third-party sources. A dagger “” denotes ratios derived from reported water consumption values when withdrawal-based ratios are not explicitly disclosed. A dash “–” indicates that the corresponding value is not reported and cannot be plausibly estimated. Approximately 94% of Hyperscale-1’s water use is potable, although the source is not explicitly identified as “municipal water.” The details of data sources are available in Appendix A.4.
@@ 3.1.2 High peaking factor
Figure 2c presents the monthly water withdrawal of a hyperscale data center in Iowa, indicating a ratio of peak monthly withdrawal to average monthly withdrawal of 4.30. It is important to note that the disclosed withdrawal volume includes a small portion of domestic water use, which typically exhibits a peaking factor in the range of 1.5 to 2.5. Consequently, the monthly peak-to-average water withdrawal ratio attributable solely to the data center cooling load is likely higher than 4.30. Applying a minimum adjustment factor of 1.5 to translate average demand to maximum day demand [92] yields an estimated daily peaking factor of at least 6.45.
We further present in Figure 2b the percentage contribution of the data center’s monthly water use relative to the total water use of the top 20 largest users served by the same water works. This analysis further highlights the pronounced seasonal variability of data centers that rely on evaporative assistance during the summer. Although the data center’s water use is relatively low during cooler months, its elevated summer water demand is sufficiently large that it becomes the largest annual water user in both 2024 and 2025 served by the water works [51].
@@
Figure 2: Water use data collected from the monthly financial reports available at [51]. (a) Monthly water use of a hyperscale data center in Iowa; (b) The ratio of the data center’s water use to the total of 20 largest water users served by same water works; (c) Monthly peaking factors (i.e., ratio of the maximum monthly water use to the average use between 2022 and 2025) of all the large water users counted by the water works.
By comparison, we also present in Figure 2c the monthly peaking factors for other large water users reported in the financial disclosures from 2022 to 2025 [51]. Some water users appear in the reports for fewer than four years; for these users, we compute the ratio of their maximum monthly water use to their average monthly use over the period in which they are reported. We can see that all other large water users exhibit substantially lower peaking factors than the data center. Note that, water user T, a large corporate campus, exhibits an abnormally high water use of approximately 15 million gallons in August 2023, exceeding more than ten times its average monthly use between 2022 and 2025. This spike is likely associated with campus expansion and infrastructure upgrades during that period, as indicated by the city’s building permit reports (May to July 2023) [95], resulting in a one time increase in water use. It does not exhibit the recurring annual pattern observed for the data center and other users. We therefore exclude this anomalous month when calculating the monthly peaking factor for user T.
@@
The water use of data centers exhibits pronounced seasonal and even daily spikes, resulting in peaking factors estimated to range from 3 to 10, depending on ambient temperature and cooling system configuration. Cooling systems that rely on evaporative assistance during the summer generally exhibit substantially higher peaking factors than conventional evaporative cooling towers. Additionally, in cooler climates and/or when server-level cooling operates at higher temperature setpoints that only need evaporative assistance for a limited number of days, the peaking factor can be even higher. For planning purposes, similar to power capacity redundancy [50, 97], data centers often request more water than their actual demand to hedge against extreme heat events and/or rapid future load increases. Finally, we note that while water storage tanks can buffer short-term fluctuations in hourly demand, they are often less effective for prolonged events such as extended or month-long heatwaves that data centers typically prepare for. As a result, data centers employing evaporative cooling generally reserve substantial daily peak water capacity allocations to hedge against these risks. Nonetheless, as water capacity is increasingly becoming a constrained societal resource (e.g., in some cases, it can take 20 years or longer to identify, develop, and bring new water supply sources into service [32]), the combination of water tanks and other water-aware computing-based solutions (Appendix F) presents a promising approach to lowering the peak water demand and strengthening public water system resilience.
@@ 4.1.2 Scenarios
While individual companies may achieve greater WUE reductions, the industry-wide adjustments applied in our study are consistent with previous observations and align with common reduction targets reported by data center operators [115, 116]. For example, a major technology company’s goal of reducing water intensity by 40% by 2030 [116], relative to its 2020 level, lies between our Moderate and Optimistic scenarios. Importantly, under the Moderate and Optimistic scenarios, the projected U.S.-wide average WUE in 2030 (Table 8) is approximately 0.43 L/kWh and 0.31 L/kWh, respectively. These values are lower than the modeled WUE projections reported in [2] and are consistent with prevailing industry trends, thereby representing plausible forward-looking estimates. In addition to the Baseline, Moderate, and Optimistic scenarios, we include two references—Reference (LBNL) and Reference (NS)—which apply WUE values based on the modeled results in [2] and [5], respectively. [^4] For our analysis, we re-calculate the total water consumption for Reference (LBNL) rather than directly using the 2024–2028 values reported in [2]. Specifically, we follow the standard annual on-site WUE definition, resulting in lower water consumption estimates than those originally reported. See Table 9 in the appendix for details.